# Low dimensional representation of multi-patient flow cytometry datasets using optimal transport for minimal residual disease detection in leukemia

### v1

@@ 3.3 Silhouette scores
We finally report in Table 2 silhouette scores to evaluate how the various embeddings of the DataML dataset displayed in Figure 4 allow to discriminate clusters. We consider the following clusters: diagnosis, normal bone marrow, positive follow-ups, and negative follow-ups. We leave out the follow-ups for which we do not have a molecular biology MRD. From the results reported in Table 2, it can be seen that the clustering obtained with OT linearization is still the best one.
Table 2: Silhouette scores for the different embeddings on DataML. For each method, we compute the score either directly on raw data or on the data projected onto the first two principal components of the PCA. The best score per row is indicated in bold.
@@ 4 Discussion
In this study, we have introduced a novel approach, via mean measure quantization and linearized OT, for low-dimensional embedding in a linear space of multi-patient FCM datasets viewed as discrete probability measures with large dimensional supports. We assessed the benefits of OT linearization over compositional data analysis and kernel mean embedding by showing that our OT-based approach yields the best silhouette scores for patients clustering. The biological relevance of our OT-based embedding has also been assessed by relating the consistency of our low-dimensional representation of multi-patient FCM datasets to the level of estimated MRD in follow-up using either molecular biology or the FlowSOW method. Interestingly, patients with positive follow-up and large MRD values tend to be clustered, in our low-dimensional representation, with data points representing FCM from diagnosis. These results suggest that using OT for statistical learning from multi-patient FCM datasets is an interesting research direction to quantify the level of MRD in the study of LAM.
@@ 5 Acknowledgments
This work benefited from financial support from the French government managed by the National Agency for Research under the France 2030 program, with the reference ANR-23-PEIA-0004.

### v2

@@ 3.3 Silhouette scores
We finally report in Table 2 silhouette scores to evaluate how the various embeddings of the DATAML-Bordeaux dataset displayed in Figure 4 allow to discriminate clusters. We consider the following clusters: diagnosis, normal bone marrow, positive MRD-BioM follow-ups, and negative MRD-BioM follow-ups. We leave out the follow-ups for which we do not have a value MRD-BioM. From the results reported in Table 2, it can be seen that the clustering obtained with OT linearization is still the best one.
Table 2: Silhouette scores for the different embeddings on DATAML-Bordeaux. For each method, we compute the score either directly on raw data or on the data projected onto the first two principal components of the PCA. The best score per row is indicated in bold.
@@ 4 Discussion
In this study, we have introduced a novel approach, via mean measure quantization and linearized OT, for low-dimensional embedding in a linear space of multi-patient FCM datasets viewed as discrete probability measures with large dimensional supports. We assessed the benefits of OT linearization over compositional data analysis and kernel mean embedding by showing that our OT-based approach yields the best silhouette scores for patients clustering. The biological relevance of our OT-based embedding has also been assessed by relating the consistency of our low-dimensional representation of multi-patient FCM datasets to the level of estimated MRD in follow-up using either molecular biology or the FlowSOW method. Interestingly, patients with positive follow-ups and large MRD values tend to be clustered, in our low-dimensional representation, with data points representing FCM from diagnosis. These results suggest that using OT for statistical learning from multi-patient FCM datasets is an interesting research direction in order to help the experts with MRD assessment. Indeed, combined with the FlowSOM method, our approach could enhance the accuracy of the relapse prognosis which is critical for making informed treatment decisions for AML patients.
@@ 5 Acknowledgments
This work benefited from financial support from the French government managed by the National Agency for Research under the France 2030 program, with the reference ANR-23-PEIA-0004. We would like to thank the data management unit of Toulouse University Hospital and the CAPTOR (Cancer Pharmacology of Toulouse Oncopole and Region) project (ANR-11-PHUC-001) for his financial support enabling e-CRF.